. In a market there are 20 buyers of a leather bag each having identical demand function Qd = 10 - 2P and 40 producers each with identical supply curve given by P= 4+Qs where Qd and Qs are quantity demanded and supplied respectively.


1
Expert's answer
2021-09-15T16:17:39-0400
"Q_S=p-4"

"10-2p=p-4"

"3p=14"

"p=4.7"

"Q=0.7"